OpenStudy (anonymous):

Rewrite the expression by factoring out (y+5) 3y^2(y+5)-5(y+5)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Use your old friend the distributive property...

OpenStudy (skullpatrol):

$$\Huge ba - ca=(b-c)a$$

OpenStudy (skullpatrol):

Let a = y +5

OpenStudy (anonymous):

The answer I have is \[\left( 3y^2 -5\right)\left( y+5 \right)\]

OpenStudy (skullpatrol):

Do you understand how they got that answer?

OpenStudy (skullpatrol):